OpenStudy (eric_d):

Simplify (sin^2 theta-2)^2-(4cos^2 theta)

OpenStudy (eric_d):

@inkyvoyd

OpenStudy (eric_d):

@ganeshie8

OpenStudy (anonymous):

a hint: It's equal to: \[\cos ^{4}\theta-2\cos ^{2}\theta + 1\] courtesy: Our Old Gandalf(Mr.Wolfram)

OpenStudy (eric_d):

hw shld i start..@Adjax

OpenStudy (anonymous):

wait a min...Do you mean \[\sin ^{2}\theta-2^{2 - 4\cos ^{2}\theta}\]

OpenStudy (anonymous):

please reconstruct the Eq...confused me a little

OpenStudy (eric_d):

nope

OpenStudy (anonymous):

can you take care to reconstruct the EQ.(EQuation)

OpenStudy (eric_d):

(sin^2 theta-2)^2-(4cos^2 theta)

OpenStudy (anonymous):

in a mathematical symbolic form (for e.g see how I posted the equation)

OpenStudy (eric_d):